Find measurement of angle A (6x+7) and measurement of angle B (3x+20) if angle A and angle B are supplementary.
Answer by jim_thompson5910(35256)

angle A and angle B are supplementary

m angle A + m angle B = 180

6x+7 + 3x+20 = 180

9x+27=180

9x=180-27

9x=153

x=(153)/(9)

x=17

Angle A: 6x+7 = 6(17)+7 = 109 degrees

Angle B: 3x+20 = 3(17)+20 = 71 degrees